# Testing Quantum Gravity with Gravitational Waves from the ringdown of binary Black Holes coalescences : A New Frontier in Fundamental Physics

### v1

@@ 6 Detectability of QNM and quantum-modified QNM frequencies
$V_{CMB}=\sum_{n}\alpha_{n}\int_{0}^{+\infty}\Big[k_{n}(\omega_{\gamma},\,t)\,|\Psi_{n}\rangle\langle\omega_{\gamma}|+\text{h.c.}\Big]\,d\omega_{\gamma}$ (6.1)
@@
$\Phi(\omega_{\gamma})=\frac{\omega_{\gamma}^{2}}{4\pi^{3}c^{2}}\frac{1}{\exp\left(\hbar\omega_{\gamma}/K_{B}T\right)-1},$ (6.2)
@@
$\sigma_{\rm abs}\sim A_{BH}=4\pi r_{S}^{2}=16\pi\left(\frac{GM}{c^{2}}\right)^{2},$ (6.3)
@@
$\Delta\omega_{n}(\omega_{\gamma},\,t)=\frac{\partial\omega_{n}}{\partial M}\frac{\hbar\,\omega_{\gamma}}{c^{2}}\simeq-\frac{\omega_{n}}{M}\frac{\hbar\,\omega_{\gamma}}{c^{2}}=-\alpha_{n}\frac{\hbar\,\omega_{\gamma}}{GM^{2}},$ (6.4)
where the QNM frequency $\omega_{n}=\alpha_{n}c^{3}/GM$ has been used. This is a semi-classical estimate, providing the order of magnitude, assuming linear response and neglecting back-reaction effects. Under this approximation for $\Delta\omega_{n}(\omega_{\gamma})$ the value of $g_{n}(\omega_{\gamma}$ in eq. 6.1 can be written as:
$g_{n}(\omega_{\gamma})\simeq\hbar\,\Delta\omega_{n}(\omega_{\gamma},\,t)\sqrt{\Phi(\omega_{\gamma})\,\sigma_{\rm abs}(\omega_{\gamma})}\,\Delta t$ (6.5)
@@
$\Gamma_{n}=2\hbar\,\gamma_{n}+\Gamma_{n}^{cont},$ (6.6)
@@
$\Gamma_{n}^{cont}=2\,\mathfrak{Im}(\Sigma_{n}(E))\sim 2\pi|\Delta\omega_{n}(\omega_{\gamma},\,t)|^{2}\,\Phi(\omega_{\gamma})\,\sigma_{\rm abs}(\omega_{\gamma})\,\Delta t^{2}.$ (6.7)
@@ 7 Quantum Gravity in the GW propagation
$E^{2}=p^{2}\left(1+\epsilon\left(\frac{E}{E_{Pl}}\right)^{\alpha}\right),$ (7.1)
@@
$\bar{c}=c\left(1+\frac{\epsilon}{2}\left(\frac{E}{E_{Pl}}\right)^{\alpha}\right),$ (7.2)
@@
$\Delta L\sim L^{1-\alpha}l_{Pl}^{\alpha},$ (7.3)
@@
$\delta h\simeq\frac{\Delta L}{L}$ (7.4)
@@
$\Delta\phi=\frac{1}{\hbar}\int{mc^{2}\,\delta g^{00}\,dt}\simeq\frac{2\pi\Delta L}{\lambda},$ (7.5)
@@
$C(\Delta\phi)=\exp\left({-\frac{1}{2}\langle(\Delta\phi)^{2}\rangle}\right)$ (7.6)
@@
$h_{**}(t,\,x)\rightarrow h_{**}C(\Delta\phi)=h_{**}(t,\,x)\exp\left({-\frac{1}{2}\langle(\Delta\phi)^{2}\rangle}\right).$ (7.7)
@@
$\mathcal{L}_{\rm CPT}\sim(k_{V})_{\kappa}\,\epsilon^{\kappa\lambda\mu\nu}\,h_{\lambda\alpha}\,\partial_{\mu}\,h_{\nu}{}^{\alpha}$. (7.8)
@@
$\omega_{\pm}^{2}=|\mathbf{k}|^{2}\pm(\kappa_{V}\cdot\mathbf{k})\,|\mathbf{k}|^{2}$ (7.9)
@@
$\Delta\Phi=\int_{0}^{L}(\omega_{+}-\omega_{-})\,dt\sim(\kappa_{V}\cdot\hat{\mathbf{k}})\,L.$ (7.10)
@@
$(\kappa_{V})_{(min)}=\frac{\Delta\Phi_{min}}{L}.$ (7.11)
@@ 8 Experimental Perspectives
$\frac{\delta h}{h}=\frac{1}{SNR}\frac{1}{\sqrt{T\,f}},$ (8.2)
@@
The evaluation of the QG impact on propagation of GWs is important for determining its effects on detector sensitivity and the possibility of observing quantum signatures in BH structure. By applying the perturbation strain as obtained in eq. 7.7, it is possible to compute the predicted correction caused by decoherence foreseen in various scenarios of QG. For typical propagation distances, the decoherence contributions predicted by the DSR scenario, the holographic model, and the random walk model are negligible. For example, in the case of the holographic model -which produces the largest contribution- for propagation distances on the order of tens or hundreds of Mpc, the ratio $\delta h/h<10^{-4}$ for decoherence effects is expected—currently beyond detection limits, but possibly accessible with future instruments. The DSR case is independent of the source distance and yields a ratio $\delta h/h\sim 10^{-70}$, while the random walk model does depend on distance but remains far below the holographic case. Decoherence effects, therefore, are relatively limited and, depending on the model considered, may only introduce a completely negligible background noise, undetectable by current instruments. The modification of the GW propagation speed can also cause a deformation of the expected spectrum, but this effect appears to be subdominant compared to other expected perturbations caused, for instance, by the presumed quantization of BHs. The only way to verify whether a presumed quantization of spacetime can affect the propagation of GWs is, therefore, to compare their propagation speed with that of other cosmic messengers that might be affected in a different way by QG.
@@ Acknowledgments
The authors would like to express their sincere gratitude to Stefano Liberati for his valuable advice, which greatly helped to improve this research work.

### v2

@@ 6 Detectability of QNM and quantum-modified QNM frequencies
$V_{CMB}=\sum_{n}\alpha_{n}\int_{0}^{+\infty}\Big[k_{n}(\omega_{\gamma},\,t)\,|\Psi_{n}\rangle\langle\omega_{\gamma}|+\text{h.c.}\Big]\,d\omega_{\gamma}$ (45)
@@
$\Phi(\omega_{\gamma})=\frac{\omega_{\gamma}^{2}}{4\pi^{3}c^{2}}\frac{1}{\exp\left(\hbar\omega_{\gamma}/K_{B}T\right)-1},$ (46)
@@
$\sigma_{\rm abs}\sim A_{BH}=4\pi r_{S}^{2}=16\pi\left(\frac{GM}{c^{2}}\right)^{2},$ (47)
@@
$\Delta\omega_{n}(\omega_{\gamma},\,t)=\frac{\partial\omega_{n}}{\partial M}\frac{\hbar\,\omega_{\gamma}}{c^{2}}\simeq-\frac{\omega_{n}}{M}\frac{\hbar\,\omega_{\gamma}}{c^{2}}=-\alpha_{n}\frac{\hbar\,\omega_{\gamma}}{GM^{2}},$ (48)
where the QNM frequency $\omega_{n}=\alpha_{n}c^{3}/GM$ has been used. This is a semi-classical estimate, providing the order of magnitude, assuming linear response and neglecting back-reaction effects. Under this approximation for $\Delta\omega_{n}(\omega_{\gamma})$ the value of $g_{n}(\omega_{\gamma}$ in eq. 45 can be written as:
$g_{n}(\omega_{\gamma})\simeq\hbar\,\Delta\omega_{n}(\omega_{\gamma},\,t)\sqrt{\Phi(\omega_{\gamma})\,\sigma_{\rm abs}(\omega_{\gamma})}\,\Delta t$ (49)
@@
$\Gamma_{n}=2\hbar\,\gamma_{n}+\Gamma_{n}^{cont},$ (50)
@@
$\Gamma_{n}^{cont}=2\,\mathfrak{Im}(\Sigma_{n}(E))\sim 2\pi|\Delta\omega_{n}(\omega_{\gamma},\,t)|^{2}\,\Phi(\omega_{\gamma})\,\sigma_{\rm abs}(\omega_{\gamma})\,\Delta t^{2}.$ (51)
@@ 7 Quantum Gravity in the GW propagation
$E^{2}=p^{2}\left(1+\epsilon\left(\frac{E}{E_{Pl}}\right)^{\alpha}\right),$ (52)
@@
$\bar{c}=c\left(1+\frac{\epsilon}{2}\left(\frac{E}{E_{Pl}}\right)^{\alpha}\right),$ (53)
@@
$\Delta L\sim L^{1-\alpha}l_{Pl}^{\alpha},$ (54)
@@
$\delta h\simeq\frac{\Delta L}{L}$ (55)
@@
$\Delta\phi=\frac{1}{\hbar}\int{mc^{2}\,\delta g^{00}\,dt}\simeq\frac{2\pi\Delta L}{\lambda},$ (56)
@@
$C(\Delta\phi)=\exp\left({-\frac{1}{2}\langle(\Delta\phi)^{2}\rangle}\right)$ (57)
@@
$h_{**}(t,\,x)\rightarrow h_{**}C(\Delta\phi)=h_{**}(t,\,x)\exp\left({-\frac{1}{2}\langle(\Delta\phi)^{2}\rangle}\right).$ (58)
@@
$\mathcal{L}_{\rm CPT}\sim(k_{V})_{\kappa}\,\epsilon^{\kappa\lambda\mu\nu}\,h_{\lambda\alpha}\,\partial_{\mu}\,h_{\nu}{}^{\alpha}$. (59)
@@
$\omega_{\pm}^{2}=|\mathbf{k}|^{2}\pm(\kappa_{V}\cdot\mathbf{k})\,|\mathbf{k}|^{2}$ (60)
@@
$\Delta\Phi=\int_{0}^{L}(\omega_{+}-\omega_{-})\,dt\sim(\kappa_{V}\cdot\hat{\mathbf{k}})\,L.$ (61)
@@
$(\kappa_{V})_{(min)}=\frac{\Delta\Phi_{min}}{L}.$ (62)
@@ 8 Experimental Perspectives
$\frac{\delta h}{h}=\frac{1}{SNR}\frac{1}{\sqrt{T\,f}},$ (64)
@@
The evaluation of the QG impact on propagation of GWs is important for determining its effects on detector sensitivity and the possibility of observing quantum signatures in BH structure. By applying the perturbation strain as obtained in eq. 58, it is possible to compute the predicted correction caused by decoherence foreseen in various scenarios of QG. For typical propagation distances, the decoherence contributions predicted by the DSR scenario, the holographic model, and the random walk model are negligible. For example, in the case of the holographic model -which produces the largest contribution- for propagation distances on the order of tens or hundreds of Mpc, the ratio $\delta h/h<10^{-4}$ for decoherence effects is expected—currently beyond detection limits, but possibly accessible with future instruments. The DSR case is independent of the source distance and yields a ratio $\delta h/h\sim 10^{-70}$, while the random walk model does depend on distance but remains far below the holographic case. Decoherence effects, therefore, are relatively limited and, depending on the model considered, may only introduce a completely negligible background noise, undetectable by current instruments. The modification of the GW propagation speed can also cause a deformation of the expected spectrum, but this effect appears to be subdominant compared to other expected perturbations caused, for instance, by the presumed quantization of BHs. The only way to verify whether a presumed quantization of spacetime can affect the propagation of GWs is, therefore, to compare their propagation speed with that of other cosmic messengers that might be affected in a different way by QG.
@@ Acknowledgments
The authors would like to express their sincere gratitude to Stefano Liberati, Vittorio Gorini and Sergio Cacciatori for their valuable advice, which greatly helped to improve this research work.